The Nutrition Facts panel on a food label lists the following information for amounts per serving: 180 Calories. There are 4 grams of total fat per serving. What percentage of the total Calories are provided by fat?

30%

25%

15%

10%

20%
Answer #1

1 gram of fat produces 9 calorie

Total calaries per serving = 180 calories

Total fat per serving = 4 grams

So total energy produced by 4 grams of fat = 4 × 9 = 36 grams

So energy provided by fat only = 36 / 180 = 0.2 × 100 = 20%

To convert in percent multiplied by 100

So the correct answer would be 20%